OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. x2 + 16 = -10x A. -2 B. -10 C. -8 D. 10 E. 8 F. 2

OpenStudy (malijhaa):

Let's solve your equation step-by-step. x2+16=−10x Step 1: Subtract -10x from both sides. x2+16−−10x=−10x−−10x x2+10x+16=0 Step 2: Factor left side of equation. (x+2)(x+8)=0 Step 3: Set factors equal to 0. x+2=0 or x+8=0 x=−2 or x=−8

OpenStudy (malijhaa):

Answer: x=−2 or x=−8

OpenStudy (anonymous):

explain?

OpenStudy (malijhaa):

Thats The Best It Could Be Explained :/ The Steps AND Answer Is There

OpenStudy (anonymous):

can u help wit more?

OpenStudy (malijhaa):

Yess

OpenStudy (anonymous):

Look at the quadratic equation below. -x2 + 4 = (3x - 1)2 What is this quadratic equation written in standard form? A. 10x2 - 6x - 3 = 0 B. 10x2 - 6x + 5 = 0 C. 8x2 - 6x + 5 = 0 D. 8x2 - 6x - 3 = 0

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

?

OpenStudy (malijhaa):

Its B. Its Very Hard To Explain

OpenStudy (anonymous):

ty <3

OpenStudy (malijhaa):

Yhur Welcome :)

OpenStudy (anonymous):

another?

OpenStudy (malijhaa):

Yup

OpenStudy (anonymous):

Look at the quadratic equation below. x2 - 2 = -2x2 + 5x What is this quadratic equation written in standard form? A. 3x2 + 5x - 2 = 0 B. x2 + 5x - 2 = 0 C. x2 - 5x - 2 = 0 D. 3x2 - 5x - 2 = 0

OpenStudy (malijhaa):

C.

OpenStudy (anonymous):

can i make new post?